OpenStudy (yamyam70):

use half angle identity for :) cos 3pi/8

OpenStudy (debbieg):

http://www.purplemath.com/modules/idents.htm Use the ID for cos(x/2) Write \(\Large \dfrac{3\pi}{8}\) as \(\Large \dfrac{1}{2}\cdot x \) , so x=? (It will be one of your unit circle angle measures, so you know the cos(x) value that you have to use in the identity) Then just plug & chug in that identity! You know this is in the first quadrant, so you know the sign of the result.

OpenStudy (yamyam70):

I don;t get the part about 1/2 * x thats where I'm getting confuseed

OpenStudy (yamyam70):

@DebbieG

OpenStudy (yamyam70):

is it 3pi over 4 ?

OpenStudy (debbieg):

lol... this is similar to the apples. :) You need to figure out an x, so that you can write the angle as \(\Large \dfrac{1}{2}\cdot x\) In other words, you need: \(\Large \dfrac{3\pi}{8}=\dfrac{1}{2}\left(???\right)\)

OpenStudy (yamyam70):

apples :D

OpenStudy (debbieg):

Right! \(\Large \dfrac{3\pi}{8}=\dfrac{1}{2}\left(\dfrac{3\pi}{4}\right)\) yes, apples... :)

OpenStudy (yamyam70):

Thank you thank you thank you! so in half angle Identity should be , cos 3pi/2 ______ ? 2

OpenStudy (debbieg):

Not quite. You want: \(\Large \cos\dfrac{3\pi}{8}=\cos\dfrac{x}{2}=\cos\dfrac{3\pi/4}{2}\) Remember, you figured out above that your \(x=\dfrac{3\pi}{4}\)

OpenStudy (yamyam70):

so we say 3pi/8 over 2 is equal to 3pi / 4 ?

OpenStudy (yamyam70):

apples..

OpenStudy (debbieg):

Because, \(\Large \dfrac{3\pi}{8}=\dfrac{1}{2}\cdot \dfrac{3\pi}{4}\)

OpenStudy (debbieg):

No, you have it backwards. NOT: "3pi/8 over 2 is equal to 3pi / 4 ? "

OpenStudy (debbieg):

But: 3pi/8 is equal to (3pi / 4) over 2

OpenStudy (yamyam70):

I see! and we get the half angle identity then?

OpenStudy (debbieg):

Remember, dividing by 2 is EXACTLY THE SAME as multiplying by 1/2, so it's easier to look at these kinds of expressions as: \(\Large \dfrac{3\pi}{8}=\dfrac{1}{2}\cdot \dfrac{3\pi}{4}\) Yes, use the ID for cos(x/2), with \(\Large x= \dfrac{3\pi}{4}\)

OpenStudy (yamyam70):

having cos 3pi / 4 over 2 = (identity formula )

OpenStudy (yamyam70):

Thanks! again @DebbieG !!!!

OpenStudy (debbieg):

You're welcome. :)

OpenStudy (yamyam70):

:)